In 1995 the United States recovered 27% of its municipal wastes through recycling, up from 17% in 1990. Let P represent the percentage recycled and t the number of years since 1990. Find a linear function P(t) that fits this data. Use this function to predict the percentage recycled in 2002.
Answer by Boreal(15235)

let 1990=0 and 1995=5
so (0, 17) and (5, 27) are points
that slope is 2
point slope formula y-y1=m(x-x1) where m is slope and (x1, y1) point
y-17=2(x)
y=2x+17
2002 is x=12
so y=41 and units are %
41%